OpenStudy (anonymous):

Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. If Laura returned the movie and paid $5.00, how many nights did she keep the movie? 14 1.5 8 6

OpenStudy (jhannybean):

5.00 = 2.00 +0.50x

OpenStudy (jhannybean):

Solve for x. \[\large \cfrac{5-2}{0.50}=x\]

OpenStudy (jhannybean):

Then check your answer. \[\large 2.00 +0.50\left(\cfrac{5-2}{0.50}\right) = 5?\]

OpenStudy (anonymous):

6

OpenStudy (jhannybean):

good job.

OpenStudy (jhannybean):

2.00 +0.50(6) = 5.00 \(\checkmark\)